OpenStudy (anonymous):

Find the vertices and foci of the hyperbola with equation x^2/9 -y^2/16 = 1 1.Vertices: (± 5, 0); Foci: (± 3, 0) 2.Vertices: (± 3, 0); Foci: (± 5, 0) 3.Vertices: (0, ± 3); Foci: (0, ±5) 4.Vertices: (0, ± 5); Foci: (0, ± 3)

OpenStudy (anonymous):

everyone hates these conic section problems i could spend all my time here doing them because no one else likes them do you know what this thing looks like? you need that before we start

OpenStudy (anonymous):

no I don't know :/

OpenStudy (anonymous):

the bigger number is under the \(y\) that means it looks like this 

OpenStudy (anonymous):

the center is \((0,0)\) so the vertices are 3 units to the left and right of that

OpenStudy (anonymous):

think of it as \[\frac{x^2}{a^2}-\frac{y^2}{b^2}=1\]oh i said something completely wrong before let me clear that up

OpenStudy (anonymous):

it looks like that because the \(x^2\) term is first! not because of the bigger number thing, that is for an ellipse

OpenStudy (anonymous):

okay i'm following.

OpenStudy (anonymous):

so if you know it looks like that, and you know the center is \((0,0)\) and you also know \(a^2=9\) making \(a=3\) you have the vertices at \((-3,0)\) and \((3,0)\)

OpenStudy (anonymous):

i.e. 3 units to the left and right that is why you need to know how it is oriented first, to know if it is left and right or up and down

OpenStudy (anonymous):

now for the foci, all you need is "c" which you find via pythagoras \[c^2=a^2+b^2\]

OpenStudy (anonymous):

so c=5^2

OpenStudy (anonymous):

i think you mean \(c^2=5^2\) so \(c=5\) right

OpenStudy (anonymous):

yeah that's what i meant :)

OpenStudy (anonymous):

and then you are basically done you know the foci are 5 units to the right and left of \((0,0)\)

OpenStudy (anonymous):

-5,0 5,0

OpenStudy (anonymous):

exactly go with 2.Vertices: (± 3, 0); Foci: (± 5, 0) again, that is why you need to know what it looks like to start, otherwise none of this is possible

OpenStudy (anonymous):

thank you so much!

OpenStudy (anonymous):

yw love them conix!